# Intrinsic spin currents in bulk noncentrosymmetric ferromagnets

### v1

@@ 1 Introduction
The present paper addresses several topics in the theory of the intrinsic spin currents in bulk noncentrosymmetric ferromagnets. The focus is on systems with the exchange splitting dominating over the strength of SO interaction. Different tools are used in the study: perturbation expansion, group theory, and computations on model as well as ab initio levels. Two particular systems are chosen for first-principles calculations, namely, NiMnSb with $C1_{b}$ structure [27] and CoMnFeSi with LiMgPdSn-type structure [28], both featured by the crystallographic point group $\bar{4}3m$. Special attention is paid to the dependence of the spin currents on direction of magnetization. The main aim is to get insight into the properties of intrinsic spin currents for a comparison with the known features of field-induced spin currents [8, 9, 10] and of the DM interaction [21].
@@ 3.1 Polarization of spin currents
However, the scaling of the projection vector $P_{\lambda}$, Eq. (3), is featured by an exponent higher than unity. This can be seen from the expansion $P_{\lambda}=P^{(0)}_{\lambda}+P^{(1)}_{\lambda}+\dots$, where the individual terms correspond to those in Eq. (6). We get not only $P^{(0)}_{\lambda}=0$ again, but also $P^{(1)}_{\lambda}=0$. The latter follows from Eqs. (3), (4), and (6), which yield
@@
$\mathrm{Tr}\{(\mathbf{m}\cdot\bm{\sigma})V_{\lambda}G_{0}(\epsilon)H^{\mathrm{SO}}G_{0}(\epsilon)\}$ $=$
$\mathrm{Tr}\{G_{0}(\epsilon)V_{\lambda}G_{0}(\epsilon)H^{\mathrm{SO}}(\mathbf{m}\cdot\bm{\sigma})\}$ $=$ $0$ (8)
@@ 3.3 Spin currents in NiMnSb and CoMnFeSi alloys
The real systems selected for ab initio study are NiMnSb with $C1_{b}$ structure [27] and CoMnFeSi with LiMgPdSn-type structure [28], both belonging to the crystallographic space group $F\bar{4}3m$ (No. 216). Their structures are derived from an fcc Bravais lattice with four sublattices, labelled $A$, $B$, $C$, and $D$, which are shifted mutually along the (111) direction of the cubic lattice. The basis vectors of these sublattices are: $A(0,0,0)$, $B(0.25,0.25,0.25)$, $C(0.5,0.5,0.5)$, and $D(0.75,0.75,0.75)$, all in units of the fcc lattice parameter. The sublattice occupation in both compounds is: Ni($A$)Mn($B$)Sb($D$) (with $C$ sublattice empty) and Co($A$)Mn($B$)Fe($C$)Si($D$). Their total magnetic moment is around $4~{}\mu_{\mathrm{B}}$ per formula unit [27, 28], indicating that both systems are in a regime of strong exchange splitting.
@@ 3.4 Dependence on magnetization direction
$T_{\kappa\lambda\mu\nu}=\sum_{\kappa^{\prime}\lambda^{\prime}\mu^{\prime}\nu^{\prime}}\alpha_{\kappa\kappa^{\prime}}\alpha_{\lambda\lambda^{\prime}}\alpha_{\mu\mu^{\prime}}\alpha_{\nu\nu^{\prime}}T_{\kappa^{\prime}\lambda^{\prime}\mu^{\prime}\nu^{\prime}},$ (10)
where $\{\alpha_{\mu\nu}\}$ are real orthogonal $3\times 3$ matrices representing the elements $\alpha$ of the point group. These conditions and the symmetry property define the shape of the tensor $T_{\kappa\lambda\mu\nu}$. The numbers of independent components of this tensor for all noncentrosymmetric point groups are listed in Table 4.
@@ Appendix A One-dimensional tight-binding model
The structure of the one-dimensional model is depicted in Fig. 5: a periodic chain along the $x$ axis comprising one magnetic site and one nonmagnetic site in the unit cell, with all sites lying in the $x-y$ plane.

### v3

@@ 1 Introduction
The present paper addresses several topics in the theory of the intrinsic spin currents in bulk noncentrosymmetric ferromagnets. The focus is on systems with the exchange splitting dominating over the strength of SO interaction. Different tools are used in the study: perturbation expansion, group theory, and computations on model as well as ab initio levels. Two particular Heusler-like systems are chosen for first-principles calculations, namely, NiMnSb with $C1_{b}$ structure [27] and CoMnFeSi with LiMgPdSn-type structure [28], both featured by the crystallographic point group $\bar{4}3m$. Special attention is paid to the dependence of the spin currents on direction of magnetization. The main aim is to get insight into the properties of intrinsic spin currents for a comparison with the known features of field-induced spin currents [8, 9, 10] and of the DM interaction [21].
@@ 3.1 Polarization of spin currents
However, the scaling of the projection vector $P_{\lambda}$, Eq. (3), is featured by an exponent higher than unity. This can be seen from the expansion $P_{\lambda}=P^{(0)}_{\lambda}+P^{(1)}_{\lambda}+\dots$, where the individual terms correspond to those in Eq. (6). We get not only $P^{(0)}_{\lambda}=0$ (needed for a linear scaling), but also $P^{(1)}_{\lambda}=0$ (leading to a quadratic scaling). The latter follows from Eqs. (3), (4), and (6), which yield
@@
$F(\epsilon)=\mathrm{Tr}\{G_{0}(\epsilon)V_{\lambda}G_{0}(\epsilon)H^{\mathrm{SO}}(\mathbf{m}\cdot\bm{\sigma})\}=0.$ (8)
@@ 3.3 Spin currents in NiMnSb and CoMnFeSi alloys
The real systems selected for ab initio study are NiMnSb with $C1_{b}$ structure [27] and CoMnFeSi with LiMgPdSn-type structure [28], both belonging to the crystallographic space group $F\bar{4}3m$ (No. 216). Their structures (related closely to the standard Heusler structure) are derived from an fcc Bravais lattice with four sublattices, labelled $A$, $B$, $C$, and $D$, which are shifted mutually along the (111) direction of the cubic lattice. The basis vectors of these sublattices are: $A(0,0,0)$, $B(0.25,0.25,0.25)$, $C(0.5,0.5,0.5)$, and $D(0.75,0.75,0.75)$, all in units of the fcc lattice parameter. The sublattice occupation in both compounds is: Ni($A$)Mn($B$)Sb($D$) (with $C$ sublattice empty) and Co($A$)Mn($B$)Fe($C$)Si($D$). Their total magnetic moment is around $4~{}\mu_{\mathrm{B}}$ per formula unit [27, 28], indicating that both systems are in a regime of strong exchange splitting.
@@ 3.4 Dependence on magnetization direction
$T_{\kappa\lambda\mu\nu}=|\alpha|\sum_{\kappa^{\prime}\lambda^{\prime}\mu^{\prime}\nu^{\prime}}\alpha_{\kappa\kappa^{\prime}}\alpha_{\lambda\lambda^{\prime}}\alpha_{\mu\mu^{\prime}}\alpha_{\nu\nu^{\prime}}T_{\kappa^{\prime}\lambda^{\prime}\mu^{\prime}\nu^{\prime}},$ (10)
where $\{\alpha_{\mu\nu}\}$ are real orthogonal $3\times 3$ matrices representing the elements $\alpha$ of the point group and $|\alpha|$ denotes determinant of $\alpha$. These conditions and the symmetry property define the shape of the tensor $T_{\kappa\lambda\mu\nu}$. The numbers of independent components of this tensor for all noncentrosymmetric point groups are listed in Table 4.
@@ Appendix A One-dimensional tight-binding model
The structure of the one-dimensional model is depicted in Fig. 5: an infinite periodic chain along the $x$ axis comprising one magnetic site and one nonmagnetic site in the unit cell, with all sites lying in the $x-y$ plane.